# Trees of Graphs as Boundaries of Hyperbolic Groups

### v1

@@ Abstract.
We characterize those 1-ended word hyperbolic groups whose Gromov boundaries are homeomorphic to trees of graphs, the family of simplest connected 1-dimensional spaces appearing among Gromov boundaries. The characterization refers to the features of the Bowditch JSJ splitting of the corresponding groups (i.e their canonical JSJ splitting over 2-ended subgroups).
@@ 1. Introduction
    to describe explicitely the class of topological spaces which appear as Gromov boundaries of word hyperbolic groups, and
@@
Concerning the above problem (1), the paper [22] provides its reduction to the case when the boundary is connected, i.e. the corresponding group is 1-ended. Some further partial insight into problem (1) is provided in [18], by showing that Gromov boundaries of hyperbolic groups belong to the class of topological spaces called Markov compacta. These are, roughly speaking, the spaces which can be realized as inverse limits of sequences of polyhedra built recursively out of an initial finite polyhedron according to some finite set of modification rules. In dimension 1, the easiest to describe reasonably rich class of Markov compacta is that of so called regular trees of graphs, as introduced in Section 6 of this paper. The main result of this paper is the complete characterization of those 1-ended hyperbolic groups $G$ whose Gromov boundaries $\partial G$ are trees of graphs (see Theorem A below).
@@
In the statement of our main result, we refer to the concept of a regular tree of graphs (as mentioned above), as well as to the more general concept of a tree of graphs, as introduced in Section 4. The statement refers also to some features of the Bowditch JSJ splitting of a group $G$, as introduced in [1], and recalled in detail in Section 3. This is a particular canonical splitting of $G$ over 2-ended subgroups. It is defined when $G$ is not cocompact Fuchsian. ($G$ is cocompact Fuchsian if it acts properly and cocompactly on the hyperbolic plane, a condition equivalent to $\partial G\cong S^{1}$.) Here we mention only few details concerning the Bowditch JSJ splitting, which are necessary for the statement of Theorem A.
@@
Technically, the Bowditch JSJ splitting of $G$ is the action of $G$ on the associated Bass-Serre tree ${T}_{G}$, called the Bowditch JSJ tree for $G$. This tree is naturally bipartitioned into black and white vertices. The black vertices are either rigid or flexible, and their stabilizers are the essential factors of the splitting (which are also called rigid and flexible, respectively). The white vertex stabilizers are the maximal 2-ended subgroups of $G$. The $G$-action on ${T}_{G}$ preserves all the above mentioned types of the vertices. We introduce the notion of a rigid cluster in $T_{G}$, as follows. We say that some two rigid vertices in $T_{G}$ are adjacent if their combinatorial distance in $T_{G}$ is equal to 2. A rigid cluster is a subtree of $T_{G}$ spanned on any equivalence class of the equivalence relation on the set of rigid vertices of $T_{G}$ induced by the adjacency. A rigid cluster factor of $G$ is the stabilizing subgroup of any rigid cluster, i.e. the subgroup consisting of all elements which map this cluster to itself. In particular, any rigid factor of $G$ is contained in some rigid cluster factor, but in general rigid cluster factors are larger than rigid factors, and they are in some sense ”combined” of some collections of rigid factors. See Subsection 3.2, and in particular Remark 3.6, for more detailed explanations, and for the description of a related splitting that we call the reduced Bowditch JSJ splitting of $G$.
@@
Recall finally that a graph is 2-connected if it is finite, connected, nonempty, not reduced to a single vertex, and contains no cutpoint.
@@
As a complement to Theorem A, we present also in the paper the description of the structure of a regular tree of graphs appearing in condition (2) of this result, given explicitely in terms of the features of the corresponding Bowditch JSJ splitting of $G$ (see Theorem 10.1, and the description of the corresponding graphical connecting system $\mathcal{R}_{G}$ given in Section 10).
One can view Theorem A as confirming that regular trees of graphs are indeed the simplest topological spaces in the complexity hierarchy among connected 1-dimensional Gromov boundaries of hyperbolic groups. Theorem A shows also that the zoo of connected 1-dimensional Gromov boundaries is quite complicated, and there is a big gap in it between trees of graphs and the boundaries of those groups which contain as their rigid factors groups whose boundary is either the Sierpinski curve or the Menger curve. The closer understanding of the species inside this gap seems to be an interesting challenge for future investigations.
@@
    $G$ has no rigid factor in its (Bowditch) JSJ splitting.
@@
We now illustrate our results with few examples (see Figures 1-4). The details concerning these examples, and other examples of a different nature, will be discussed and presented in Section 14. In each of the examples below the 1-ended hyperbolic group $G_{i}$ is the fundamental group of the corresponding ”graph of surfaces” $M_{i}$. Recall also that the Bowditch JSJ splitting of a group $G$ (as well as its reduced Bowditch JSJ splitting) can be expressed by means of the related graph of groups decomposition $\mathcal{G}$ of $G$, and this is the approach that we follow below. In all four discussed examples $G_{i}$, $i\in\{1,2,3,4\}$, the edge groups of the corresponding graph of groups decompositions $\mathcal{G}_{i}$ coincide with the corresponding incident white vertex groups, so we skip the indication of the form of these edge groups at the figures.
@@
Consider the group $G_{1}=\pi_{1}M_{1}$, presented at Figure 1. The Bowditch JSJ graph of groups decomposition of $G_{1}$, denoted $\mathcal{G}_{1}$, is presented at the right part of this figure. Since all factors of this decomposition are easily seen to be flexible (because they are the conjugates of the surface subgroups $\pi_{1}F_{i}$ equipped with peripheral subgroups corresponding to the boundary curves), condition (3) of Theorem B holds for $G_{1}$, and hence the boundary $\partial G_{1}$ is homeomorphic to some regular tree of thick $\theta$-graphs.
Figure 2 shows an example in which the corresponding group $G_{2}=\pi_{1}M_{2}$ contains, up to conjugacy, precisely one rigid factor in its Bowditch JSJ splitting. This factor corresponds to the subgroup $\pi_{1}H_{0}<\pi_{1}M_{2}$, and it is obviously isomorphic to the free group of rank 2. Moreover, it is not hard to observe that the rigid clusters in the Bowditsch’s JSJ tree of $G_{2}$ coincide with the rigid vertices, and hence the rigid cluster factors coincide with the rigid factors (actually, in this case the reduced Bowditch JSJ splitting coincides with the ordinary Bowditch JSJ splitting). It follows that the group $G_{2}$ satisfies condition (3) of Theorem A, and hence its Gromov boundary $\partial G_{2}$ is homeomorphic to some regular tree of 2-connected graphs. In Section 14 we will give a complete description of this regular tree of graphs, in terms of the so called graphical connecting system $\mathcal{R}_{G_{2}}$ associated to $G_{2}$ in the way presented in Section 10.
@@
The group $G_{3}=\pi_{1}M_{3}$ presented at Figure 3 contains, up to conjugacy, precisely two rigid factors in its Bowditch JSJ splitting, and they correspond to the subgroups $\pi_{1}H_{i}<\pi_{1}M_{3}$, $i=1,2$. It can be deduced from the corresponding graph of groups decomposition $\mathcal{G}_{3}$ of $G_{3}$ that the rigid clusters are in this case nontrivial, and that the rigid cluster factors of $G_{3}$ are the conjugates of the subgroup $\pi_{1}(H_{1}\cup H_{2})$. (The corresponding reduced Bowditch JSJ graph of groups decomposition of $G_{3}$ is presented as $\mathcal{G}^{r}_{3}$ at the same figure.) Since obviously these rigid cluster factors are free, condition (3) of Theorem A is again satisfied, and thus the boundary $\partial G_{3}$ is also homeomorphic to some regular tree of 2-connected graphs.
@@
Our final example, presented at Figure 4, shows some 1-ended hyperbolic group $G_{4}=\pi_{1}M_{4}$ whose boundary $\partial G_{4}$ is not a tree of graphs, even though its all rigid factors are free. The rigid factors of $G_{4}$ correspond to the subgroups $\pi_{1}H_{i}<\pi_{1}M_{4}$, for $i=1,2$, and the rigid cluster factors correspond to the subgroup $\pi_{1}(H_{1}\cup H_{2})$. Since that latter group is the fundamental group of a closed surface, it is not virtually free, and hence condition (3) of Theorem A is not fulfilled. As a consequence, the boundary $\partial G_{4}$ is not homeomorphic to any tree of graphs.
@@
The paper is organized as follows. The first part of the paper (Sections 2–5) focuses on the proofs of the implications (1$\implies$(3) in Theorem A, and (1$\implies$(3) in Theorem B. The second part of the paper (Sections 6–13) deals with the proofs of the implications (3$\implies$(2) in Theorem A, and (3$\implies$(2) in Theorem B. The remaining implications ((2$\implies$(1) in Theorem A, and (2$\implies$(1) in Theorem B) necessary to get the full proofs of Theorems A and B are obvious.
Our exposition in both parts of the paper consists of many preparations for the proofs, including the complete descriptions and explanations of the terms appearing in the statements of our two main results. More precisely, as far as the first part, the organization is as follows. In Section 2 we recall the construction called the limit of a tree system of metric compacta. It is used later, in Section 4, to describe trees of graphs (and trees of $\theta$-graphs), and it is also used throughout the paper, as the main technical ingredient in the arguments. In Section 3 we recall the concept of the canonical (Bowditch) JSJ splitting of a 1-ended hyperbolic group, and we describe related concepts of the reduced Bowditch JSJ splitting, and of rigid cluster factors of this reduced splitting. In Section 5 we study some configurations of local cutpoints in trees of graphs (and trees of $\theta$-graphs), and we deduce the proofs of the implications (1$\implies$(3) in Theorem A, and (1$\implies$(3) in Theorem B.
As far as the second part of the paper, the organization is as follows. In Section 6 we introduce regular trees of graphs. Sections 7–9 provide various technicalities and terminology which allow to state in Section 10 (as Theorem 10.1) the result which describes explicitely the form of the regular trees of graphs which appear in our main results. In particular, we discuss in Section 9 some graphs (the so called Whitehead graphs, and their extended variants) which are naturally associated to those rigid factors (or rather rigid cluster factors) of the JSJ splittings which are virtually free. Sections 10–12 contain some further technical preparations, and in Section 13 we present the proofs of those parts of Theorems A and B, which identify the Gromov boundaries of the corresponding groups as explicitely described regular trees of graphs.
Finally, in Section 14, we discuss the examples presented above in this introduction, as well as some further examples, of a different sort, namely some right-angled Coxeter groups.
@@ 2. Tree systems and their limits
In this preparatory section we present (or rather recall) the construction called limit of a tree system of spaces (or, less formally, a tree of spaces). This construction is a slight extension of that presented in Section 1 of [23] (see also [19]). We will use this construction in the next sections to describe and study trees of graphs, the class of spaces on which we focus in this paper.
@@ 2.1. Some terminology and notation concerning trees
For any $v\in V_{T}$, we denote by $N^{T}_{v}$, or simply by $N_{v}$, the set of all vertices of $T$ adjacent to $v$. We denote by $d(v)=d_{T}(v)$ the degree of a vertex $v$ in $T$, i.e. the cardinality of the set $N^{T}_{v}$.
We denote by $\partial T$ the set of ends of $T$, i.e. the set of equivalence classes of rays in $T$ with respect to the relation of coincidence except possibly at some finite initial parts. We denote the end determined by a ray $\rho$ as $[\rho]$.
@@
Note that $N_{S}$ consists solely of white vertices of $T$. Note also that, in case when $S$ is reduced to a single black vertex $t$, this notation agrees with the earlier introduced notation for the set $N_{t}$.
@@ 2.2. Spaces with peripherals and tree systems
Say that a graph $\Gamma$ is essential if it is nonempty and contains no isolated vertex. (We allow that $\Gamma$ contains loop edges and multiple edges.) Given a finite essential graph $\Gamma$, we describe the associated space with peripherals called the punctured graph $\Gamma$, denoted $\Gamma^{\circ}$, as follows. Denote by $|\Gamma|$ the realization of $\Gamma$, i.e. the underlying topological space of $\Gamma$, which is a compact metric space. For each point $x\in|\Gamma|$ consider a normal neighbourhood of $x$ in $\Gamma$, i.e. a connected closed neighbourhood $B(x)$ of $x$ whose intersection with the vertex set of $\Gamma$ either coincides with $\{x\}$ (when $x$ is a vertex of $\Gamma$), or is empty. Denote by $\operatorname{int}B(x)$ and $\operatorname{bd}B(x)$ the interior and the boundary (in $|\Gamma|$) of a normal neighbourhood $B(x)$, respectively. A standard dense family of normal neighbourhoods in $\Gamma$ is any family of pairwise disjoint normal neighbourhoods in $\Gamma$ which contains normal neighbourhoods of all vertices of $\Gamma$, and whose union is dense in $|\Gamma|$. Note that each such family is automatically countable and null. Moreover, such family is unique up to a homeomorphism of the ambient space $|\Gamma|$ which is identical on the vertex set of $\Gamma$ and which maps all edges of $\Gamma$ to themselves.
@@
    $T$ is a bipartite countable tree (called the underlying tree of $\Theta$), with black and white vertices of the partition, such that the degree $d(u)$ of each white vertex $u$ is finite;
@@
    is a space with peripherals, i.e. the family $\Omega_{t}:=\{\varphi_{[t,u]}(\Sigma_{u}):u\in N_{t}\}$ is null and consists of pairwise disjoint sets; pairs $(K_{t},\Omega_{t})$ are called the constituent spaces with peripherals of $\Theta$.

### v2

@@ Abstract.
We characterize those 1-ended word hyperbolic groups whose Gromov boundaries are homeomorphic to trees of graphs (i.e. to inverse limits of graphs that have particularly simple finitary descriptions). These are groups with the simplest connected Gromov boundaries of topological dimension 1. The characterization is expressed in terms of algebraic properties of the Bowditch JSJ splitting of the corresponding groups (i.e. the canonical JSJ splitting over 2-ended subgroups).
Key words and phrases:
hyperbolic group, Gromov boundary, trees of graph, JSJ splitting
2010 Mathematics Subject Classification:
20F65, 20F67
@@ 1. Introduction
    to describe explicitly the class of topological spaces which appear as Gromov boundaries of word hyperbolic groups, and
@@
Concerning the above problem (1), the paper [Ś16] provides its reduction to the case when the boundary is connected, i.e., the corresponding group is 1-ended. Some further partial insight into problem (1) is provided in [Paw15], by showing that Gromov boundaries of hyperbolic groups belong to the class of topological spaces called Markov compacta. These are, roughly speaking, the spaces which can be realized as inverse limits of sequences of polyhedra built recursively out of an initial finite polyhedron according to some finite set of modification (or substitution) rules. In dimension 1, the easiest to describe reasonably rich class of Markov compacta is that of so called regular trees of graphs, as introduced in Section 6 of this paper. The main result of this paper is the complete characterization of those 1-ended hyperbolic groups $G$ whose Gromov boundaries $\partial G$ are trees of graphs (see Theorem A below).
@@
In the statement of our main result, we refer to the concept of a regular tree of graphs (as mentioned above), as well as to the more general concept of a tree of graphs, as introduced in Section 4. The statement refers also to some features of the Bowditch JSJ splitting of a group $G$, as introduced in [Bow98], and recalled in detail in Section 3. This splitting is a canonical splitting of a word hyperbolic group $G$ over 2-ended subgroups. It is defined when $G$ is not cocompact Fuchsian. ($G$ is cocompact Fuchsian if it acts properly and cocompactly on the hyperbolic plane, a condition equivalent to $\partial G\cong S^{1}$.) Here we mention just a few details concerning the Bowditch JSJ splitting, which are necessary for the statement of Theorem A.
@@
Technically, the Bowditch JSJ splitting of $G$ is the action of $G$ on the associated Bass-Serre tree ${T}_{G}$, called the Bowditch JSJ tree for $G$. This tree is naturally bipartitioned into black and white vertices. The black vertices are either rigid or flexible, and their stabilizers are the essential factors of the splitting (which are also called rigid and flexible, respectively). The white vertex stabilizers are maximal 2-ended subgroups of $G$. The $G$-action on ${T}_{G}$ preserves all the above mentioned types of vertices. We introduce the notion of a rigid cluster in $T_{G}$, as follows. We say that two rigid vertices in $T_{G}$ are adjacent if their combinatorial distance in $T_{G}$ is equal to 2. A rigid cluster is a subtree of $T_{G}$ spanned on any equivalence class of the equivalence relation on the set of rigid vertices of $T_{G}$ induced by adjacency. A rigid cluster factor of $G$ is the stabilizing subgroup of any rigid cluster, i.e., the subgroup consisting of all elements which map this cluster to itself. In particular, any rigid factor of $G$ is contained in some rigid cluster factor, but in general rigid cluster factors are larger than rigid factors, and they are in some sense “combined” from collections of rigid factors. See Section 3.2, and in particular Remark 3.6, for more detailed explanations, and for the description of a related splitting that we call the reduced Bowditch JSJ splitting of $G$.
@@
Recall finally that a graph is 2-connected if it is finite, connected, nonempty, not equal to a single vertex, and cutpoint-free.
@@
As a complement to Theorem A, we present also in the paper the description of the structure of a regular tree of graphs appearing in condition (2) of this result, given explicitly in terms of the features of the corresponding Bowditch JSJ splitting of $G$ (see Theorem 10.1, and the description of the corresponding graphical connecting system $\mathcal{R}_{G}$ given in Section 10). In fact, as we show in Remark 15.1, an explicit description of $\partial G$ as a regular tree of graphs can be similarly obtained from a splitting of $G$ coming from a much more general class of splittings over $2$-ended subgroups with maximal $2$-ended and virtually free vertex groups.
One can view Theorem A as confirming that regular trees of graphs are indeed the simplest topological spaces in the complexity hierarchy among connected 1-dimensional Gromov boundaries of hyperbolic groups. Theorem A shows also that the zoo of connected 1-dimensional Gromov boundaries is quite complicated, and there is a big gap in it between trees of graphs and the boundaries of those groups which contain as their rigid factors groups whose boundary is either the Sierpinski curve or the Menger curve (see Remark 15.6 for more detailed discussion of this issue). The closer understanding of the species inside this gap seems to be an interesting challenge for future investigations.
@@
    $G$ has no rigid factor in its Bowditch JSJ splitting.
@@
We now illustrate our results with few examples (see Figures 1–4). The details concerning these examples, and other examples of a different nature, will be discussed and presented in Section 14. In each of the examples below, the 1-ended hyperbolic group $G_{i}$ is the fundamental group of the corresponding “graph of surfaces” $M_{i}$. Recall also that the Bowditch JSJ splitting of a group $G$ (as well as its reduced Bowditch JSJ splitting) can be expressed by means of the related graph of groups decomposition $\mathcal{G}$ of $G$, and this is the approach that we follow below. In all four discussed examples $G_{i}$, $i\in\{1,2,3,4\}$, the edge groups of the corresponding graph of groups decompositions $\mathcal{G}_{i}$ coincide with the corresponding incident white vertex groups, so we skip the indication of the form of these edge groups in the figures.
@@
Consider the group $G_{1}=\pi_{1}M_{1}$, presented in Figure 1. The Bowditch JSJ graph of groups decomposition of $G_{1}$, denoted $\mathcal{G}_{1}$, is presented in the right part of this figure. Since all factors of this decomposition are easily seen to be flexible (because they are the conjugates of the surface subgroups $\pi_{1}F_{i}$ equipped with peripheral subgroups corresponding to the boundary curves), condition (3) of Theorem B holds for $G_{1}$, and hence the boundary $\partial G_{1}$ is homeomorphic to some regular tree of thick $\theta$-graphs.
Figure 2 shows an example in which the corresponding group $G_{2}=\pi_{1}M_{2}$ contains, up to conjugacy, precisely one rigid factor in its Bowditch JSJ splitting. This factor corresponds to the subgroup $\pi_{1}H_{0}<\pi_{1}M_{2}$, and it is obviously isomorphic to the free group of rank 2. Moreover, it is not hard to observe that the rigid clusters in the Bowditch JSJ tree of $G_{2}$ coincide with the rigid vertices, and hence the rigid cluster factors coincide with the rigid factors (actually, in this case the reduced Bowditch JSJ splitting coincides with the ordinary Bowditch JSJ splitting). It follows that the group $G_{2}$ satisfies condition (3) of Theorem A, and hence its Gromov boundary $\partial G_{2}$ is homeomorphic to some regular tree of 2-connected graphs. In Section 14 we will give an explicit description of this regular tree of graphs, in terms of the so called graphical connecting system $\mathcal{R}_{G_{2}}$ associated to $G_{2}$ in the way presented in Section 10.
@@
The group $G_{3}=\pi_{1}M_{3}$ presented in Figure 3 contains, up to conjugacy, precisely two rigid factors in its Bowditch JSJ splitting, and they correspond to the subgroups $\pi_{1}H_{i}<\pi_{1}M_{3}$, $i=1,2$. It can be deduced from the corresponding graph of groups decomposition $\mathcal{G}_{3}$ of $G_{3}$ that the rigid clusters are in this case nontrivial, and that the rigid cluster factors of $G_{3}$ are the conjugates of the subgroup $\pi_{1}(H_{1}\cup H_{2})$. (The corresponding reduced Bowditch JSJ graph of groups decomposition of $G_{3}$ is presented as $\mathcal{G}^{r}_{3}$ in the same figure.) Since obviously these rigid cluster factors are free, condition (3) of Theorem A is again satisfied, and thus the boundary $\partial G_{3}$ is also homeomorphic to some regular tree of 2-connected graphs.
@@
Our final example, presented in Figure 4, shows some 1-ended hyperbolic group $G_{4}=\pi_{1}M_{4}$ whose boundary $\partial G_{4}$ is not a tree of graphs, even though its all rigid factors are free. The rigid factors of $G_{4}$ correspond to the subgroups $\pi_{1}H_{i}<\pi_{1}M_{4}$, for $i=1,2$, and the rigid cluster factors correspond to the subgroup $\pi_{1}(H_{1}\cup H_{2})$. Since that latter group is the fundamental group of a closed surface, it is not virtually free, and hence condition (3) of Theorem A is not fulfilled. As a consequence, the boundary $\partial G_{4}$ is not homeomorphic to any tree of graphs.
@@
The paper is organized as follows. The first part of the paper (Sections 2–5) focuses on the proofs of the implications (1$\Rightarrow$(3) in Theorem A, and (1$\Rightarrow$(3) in Theorem B. The second part of the paper (Sections 6–13) deals with the proofs of the implications (3$\Rightarrow$(2) in Theorem A, and (3$\Rightarrow$(2) in Theorem B. The remaining implications ((2$\Rightarrow$(1) in Theorem A, and (2$\Rightarrow$(1) in Theorem B) necessary to get the full proofs of Theorems A and B are obvious.
Our exposition in both parts of the paper consists of many preparations for the proofs, including the complete descriptions and explanations of the terms appearing in the statements of our two main results. More precisely, in the first part, the organization is as follows. In Section 2 we recall the construction called the limit of a tree system of metric compacta. It is used later, in Section 4, to describe trees of graphs (and trees of $\theta$-graphs), and it is also used throughout the paper, as the main technical ingredient in the arguments. In Section 3.1 we recall the concept of the canonical (Bowditch) JSJ splitting of a 1-ended hyperbolic group, and we describe related concepts of the reduced Bowditch JSJ splitting, and of rigid cluster factors of this reduced splitting. In Section 5 we study some configurations of local cutpoints in trees of graphs (and trees of $\theta$-graphs), and we deduce the proofs of the implications (1$\Rightarrow$(3) in Theorem A, and (1$\Rightarrow$(3) in Theorem B.
In the second part of the paper, the organization is as follows. In Section 6 we introduce regular trees of graphs. Sections 7–9 provide various technicalities and terminology which allow to state in Section 10 (as Theorem 10.1) the result which describes explicitly the form of the regular trees of graphs which appear in our main results. In particular, we discuss in Section 9 some graphs (the so called Whitehead graphs, and their extended variants) which are naturally associated to the virtually free rigid factors (or rather rigid cluster factors) of the JSJ splittings. Sections 10–12 contain some further technical preparations, and in Section 13 we present the proofs of those parts of Theorems A and B, which identify the Gromov boundaries of the corresponding groups as explicitly described regular trees of graphs.
In Section 14, we discuss the examples presented above in this introduction, as well as some further examples, of a different sort, namely some right-angled Coxeter groups.
In Section 15, we make several final comments, including explaining how the description of $\partial G$ as a regular tree of graphs in terms of the Bowditch JSJ splitting can be generalized to a much more general class of splittings.
1.3. Acknowledgments.
The second author expresses his thanks to Jason Manning for directing his attention to Whitehead graphs as objects which could potentially play a role in solving the problem studied in this paper (which indeed turned out to be the case).
@@ 2. Tree systems and their limits
In this preparatory section we present (or rather recall) the construction called limit of a tree system of spaces (or, less formally, a tree of spaces). This construction is a slight extension of that presented in Section 1 of [Ś20] (see also [PZ16]). We will use this construction in the next sections to describe and study trees of graphs, the class of spaces on which we focus in this paper.
@@ 2.1. Some terminology and notation concerning trees
For any $v\in V_{T}$, we denote by $N^{T}_{v}$, or simply by $N_{v}$, the set of all vertices of $T$ adjacent to $v$. We denote by $d(v)=d_{T}(v)$ the degree of a vertex $v$ in $T$, i.e., the cardinality of the set $N^{T}_{v}$.
We denote by $\partial T$ the set of ends of $T$, i.e., the set of equivalence classes of rays in $T$ with respect to the relation of coincidence except possibly at some finite initial parts. We denote the end determined by a ray $\rho$ as $[\rho]$.
@@
Note that $N_{S}$ consists solely of white vertices of $T$. Note also that, in case when $S$ is equal to a single black vertex $t$, this notation agrees with the earlier introduced notation for the set $N_{t}$.
@@ 2.2. Spaces with peripherals and tree systems
Say that a graph $\Gamma$ is essential if it is nonempty and contains no isolated vertex. (We allow that $\Gamma$ contains loop edges and multiple edges.) Given a finite essential graph $\Gamma$, we describe the associated space with peripherals called the punctured graph $\Gamma$, denoted $\Gamma^{\circ}$, as follows. Denote by $|\Gamma|$ the realization of $\Gamma$, i.e., the underlying topological space of $\Gamma$, which is a compact metric space. For each point $x\in|\Gamma|$ consider a normal neighbourhood of $x$ in $\Gamma$, i.e., a connected closed neighbourhood $B(x)$ of $x$ whose intersection with the vertex set of $\Gamma$ either coincides with $\{x\}$ (when $x$ is a vertex of $\Gamma$), or is empty. Denote by $\operatorname{int}B(x)$ and $\operatorname{bd}B(x)$ the interior and the boundary (in $|\Gamma|$) of a normal neighbourhood $B(x)$, respectively. A standard dense family of normal neighbourhoods in $\Gamma$ is any family of pairwise disjoint normal neighbourhoods in $\Gamma$ which contains normal neighbourhoods of all vertices of $\Gamma$, and whose union is dense in $|\Gamma|$. Note that each such family is automatically countable and null. Moreover, such family is unique up to a homeomorphism of the ambient space $|\Gamma|$ which is identity on the vertex set of $\Gamma$ and which maps all edges of $\Gamma$ to themselves.
@@
    $T$ is a countable tree (called the underlying tree of $\Theta$), equipped with a proper colouring of its vertices in black and white (i.e. each edge has one of its vertices colored black and the other colored white), such that the degree $d(u)$ of each white vertex $u$ is finite;
@@
    is a space with peripherals, i.e., the family $\Omega_{t}:=\{\varphi_{[t,u]}(\Sigma_{u}):u\in N_{t}\}$ is null and consists of pairwise disjoint sets; pairs $(K_{t},\Omega_{t})$ are called the constituent spaces with peripherals of $\Theta$.